lucy has a rectangular plot of land with a length of 20 feet and a width of 10 feet. She wants to design a flower garden in the shape of a circle with two semicircles at each end of the center circle. She will fill in the remaining area with wood chips. If one bag of wood chips covers 5 square feet, how many bags must she buy to fill the shaded area?
Answer by mananth(16946)

Basically she has two half circles and and on circle at the centre
which means 2 full circles
The diameter of the circles will be the width of the plot which is 10 feet
so radius of the circles will be 5 ft
Area of circle = pir^2
= pi * 5^2
= 25 pi
two circles
so 50 pi is the area of the garden
=157 sq.ft
area of rectangle = 20 * 10 = 300 sq.ft
area to be covered with chips = 200 -157= 43 sq. ft
5 sq. ft ----------1 bag
43 sq. ft ----------- 43/5 = 8.6 bags
= 9 bags